For the system methane and ethane for equimolar mixture at temperature 100°C and pressure 35bar, estimate the fugacity of gases
through application of the Generlized Correlations.
Answer #1

We will first find the pseudo critical properties of the mixture to find the reduced temperature and pressure.

Critical temperature of methane is, T_{c}^{m}=190.6K

Critical pressure of methane is Pm 46bar

Critical temperature of ethane is, T_{c}^{e}=305.3K

Critical pressure of ethane is P_{c}^{e}=49bar

Now the pseudo critical temperature of the mxture will be, T_{c}'=0.5*190.6+0.5*305.3=247.95K

pseudo critical pressure of the mxture will be, P_{c}'=0.5*46+0.5*49=47.5bar

Now the given temperature is T=100^{0}C=373K and pressure is P=35bar

Now the reduced temperature is calculated as, T_{r}=\frac{T}{T_{c}'}=\frac{373}{247.95}=1.5

The reduced pressure is calculated as, P_{r}=\frac{P}{P_{c}'}=\frac{35}{47.5}=0.74

Now we can calculate the fugacity using the formulae,

ln\frac{f}{P}=\int_{0}^{P_{r}}\frac{Z-1}{P_{r}}dP_{r}

We can calculate the integral analytically using simpson's 1/3 rd rule at constant temperature by taking compressibility factors from the isotherms on the generalised compressibilty chart.

TK 500 3.00 2009 1.00 -1.60 1 0.90 o -1.40 1.5 1.30 0.80 0.85 120 15 10.90 10 0.90 1.00 0.80 0.95 PV 0.70 Z= RT 0.60 17:28 1.

From the generalised compressibility chart we will take points equidistant P_{r} values from 0 to 0.74 and we will find the corresponding .

The Simpson's 1/3rd rule is given as,

b a f(x)dx = $[F(xo)+4{f(x, J+f(x3) +...+ f(x-1)}+...] ...+2{f(x2)+ f(xx)+...+ f(x,-2)}+ f(x,)}] f(x)+4 Ef(x;)+2 Ef(x;)+F(x,

Here f(x)=\frac{Z-1}{P_{r}} and dx=dP_{r}

The limits are a=0 and b=P_{r}=0.74

We will choose an interval of h=0.02 , so there will be a total of n=37 points

We will do the rest of the calculation in excel,

0905 TL 5.00 -3.00 2.00 1 880 2009 1.60 0.9 1.46 2.20 3.00 60 -0.80 2.00 80 60 1.30 1.50 40 1.30 $120 1.10 0.8 0.85 1.20 15 1x 1 f(x) 이 1 0 0.995 | 1 | 0.02 0.995 | 2 10.04 0.995 | 3 10.06 0.995 15 | 0.1 | 0.99 | 0.12 | 0.99 | 7 8 10.14 10.16 | 0.99

We will get the value of integral as 0.686

So,

ln\frac{f}{P}=0.686

\Rightarrowf=69.5bar

So the fugacity of equimolar ethane methane mixture is 69.5 bar